Fermi National Laboratory


Accelerator Update for January 12, 2005 - January 14, 2005

Wednesday January 12
The day shift began with Operations monitoring store 3918 and the Recycler stash, and with MiniBooNE and Meson MCenter and MTest experiments taking beam.

MCenter stopped using beam at 9:19 AM.

Operations terminated store 3918 at 5:13 PM and then put the Tevatron (TeV) into shot setup at 5:37 PM.

Operations reported at 8:07 PM that the shot was on hold due to problems with the TeV BLMs. The protons had already been loaded. After rebooting several systems, operators aborted the protons and waited for expert help.

Controls experts found a software problem at 10:29 PM, which resolved the BLM problem.

Operations established store 3920 at 11:52 PM with an initial luminosity of 75.48E30. Pbar resumed stacking.

Thursday January 13
The midnight shift began with Operations monitoring store 3920, with Pbar stacking, and with MiniBooNE taking beam.

At 12:54 AM, a Collider Detector at Fermilab (CDF) experimenter called up to complain that a beam problem had caused some SVX crates to trip off. This shut down their experiment.

Operations established beam to MCenter at 1:38 AM.

At 2:07 AM, a Main Injector (MI) quadrupole (M:Q103) tripped off and wouldn't reset. Operators called in an expert.

A Booster power supply (WAPS) tripped off at 3:07 AM due to a problem with an RF station (BFR18) tripping off. This happened several times throughout the shift.

An expert came in and had M:Q103 repaired by 6:17 AM. Pbar resumed stacking.

At 6:46 AM, a Pbar power supply (PMAG) tripped due to a stuck LCW flow switch. Operators contacted an expert, who had the supply reset by 7:53 AM. Pbar resumed stacking.

At 11:43 AM, BRF18 tripped off. It reset okay.

Operations established beam to MCenter at 12:05 PM.

Operations turned off the Booster beam at 1:28 PM for tunnel work.

A Preaccelerator expert accessed the I- Source at 2:11 PM. Their access ended 30 minutes later.

Operators prepared MI for an access at 2:17 PM.

Booster and Linac work finished and beam resumed at 4:28 PM.

MiniBooNE resumed taking beam at 5:01 PM.

At 5:40 PM, Operations reported having trouble extracting beam to MiniBooNE due to a kicker (MKS05) failure. Experts investigated. During the investigation, operators lowered intensity and slowly resumed sending beam to MiniBooNE.

A Booster expert reported that he had replaced a bad cable on MKS05. He had the kicker running by 7:28 PM.

Operations contacted an expert at 7:49 PM to investigate problems with MI's MI-60 lower power supply.

At 9:48 PM, the Linac 750 keV Line Buncher tripped off.

At 10:02 PM, Operations received an alarm for MI-60 North. The power supply had been worked on and returned to service, but after ten ramps the VCB cubicle in the yard suffered an internal flashover and a fire, but the fire was short lived. The Fire Department verified the fire was out. Operators racked out its pulsed power, the transfer supplies at MI30, and bypassed the power supply. Experts were called and they began a preliminary investigation without disturbing the scene.

Operation made MI ready for an access at 10:50 PM.

Friday January 14
The midnight shift began with Operations monitoring store 3920, and with MiniBooNE taking beam. MI was off due to the MI-60 power supply/VCB problem. Operations monitored the Pbar stack. MCenter and MTest were also off.

At 6:46 AM, Operations reported that the MiniBooNE horn tripped off. It reset okay.

The Plan for Friday and the Weekend
The plan for today is to keep store 3920 in as long as possible. If the store ends for some reason, the Run Coordinator may allow machine accesses. Experts are deciding the best way to investigate the MI-60 VCB and to see if we can operate MI without the MI-60 power supply. If a safe procedure can be developed, Operations will establish a new store with the existing stack. However, at this time there is no estimate as to when this might happen or when repairs to the supply or VCB might begin.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/14/2005   email Fermilab